Two types of chemical solutions treams, A and B, were tested for their alkalinity levels. Analysis of the 6 samples taken from solution A showed that the mean alkalinity level is 7.52 with a standard deviation of 0.024 while the 5 samples taken at Solution B showed an alkalinity level of 7.49 with a standard deviation of 0.032. Using a (a) 0.05 and (b) 0.01 significance level, determine whether the streams have different alkalinity levels. Use F-distribution.
1
Expert's answer
2020-05-25T21:06:02-0400

The following null and alternative hypotheses need to be tested:

"H_0:\sigma_1^2=\sigma_2^2"

"H_1:\sigma_1^2\not=\sigma_2^2"

This corresponds to a two-tailed test, for which a F-test for two population variances needs to be used.

Based on the information provided, the significance level is "\alpha=0.05," and the the rejection region for this two-tailed test is "R=\{F:F<0.135\ or\ F>9.364\}."

The F-statistic is computed as follows:


"F={s_1^2 \over s_2^2}={0.000576 \over 0.001024}=0.5625"

Since from the sample information we get that "F_L=0.135<F=0.5625<9.364=F_U," it is then concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population variance "\sigma_1^2" is different than the population variance "\sigma_2^2," at the "\alpha=0.05" significance level.


Based on the information provided, the significance level is "\alpha=0.01," and the the rejection region for this two-tailed test is "R=\{F:F<0.064\ or\ F>22.456\}."

The F-statistic is computed as follows:


"F={s_1^2 \over s_2^2}={0.000576 \over 0.001024}=0.5625"

Since from the sample information we get that "F_L=0.064<F=0.5625<22.456=F_U," it is then concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population variance "\sigma_1^2" is different than the population variance "\sigma_2^2," at the "\alpha=0.01" significance level.


The following null and alternative hypotheses need to be tested:

"H_0:\mu_1=\mu_2"

"H_1:\mu_1\not=\mu_2"

a) This corresponds to a two-tailed test, for which a t-test for two population means, with two independent samples, with unknown population standard deviations will be used.

Based on the information provided, the significance level is "\alpha=0.05," and the degrees of freedom are "df=6+5-2=9." In fact, the degrees of freedom are computed as follows, assuming that the population variances are equal.

Hence, it is found that the critical value for this two-tailed test is "t_c=2.262," for "\alpha=0.05" and "df=9."

The rejection region for this two-tailed test is "R=\{t:|t|>2.262\}."

Since it is assumed that the population variances are equal, the t-statistic is computed as follows:


"t={\bar{X_1}-\bar{X_2}\over \sqrt{\dfrac{(n_1-1)s_1^2+(n_2-1)s_2^2}{n_1+n_2-2}(\dfrac{1}{n_1}+\dfrac{1}{n_2})}}="

"={7.52-7.49\over \sqrt{\dfrac{(6-1)(0.024)^2+(5-1)(0.032)^2}{6+5-2}(\dfrac{1}{6}+\dfrac{1}{5})}}="

"=1.78"

Since it is observed that "|t|=1.78<2.262=t_c," it is then concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that the population mean "\mu_1" is different than "\mu_2," ​at the 0.05 significance level.

Using the P-value approach: The p-value is "p=0.1089," and since "p=0.1089>0.05," it is concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that the population mean "\mu_1" is different than "\mu_2," at the 0.05 significance level.


b) Based on the information provided, the significance level is "\alpha=0.01," and the degrees of freedom are "df=6+5-2=9." In fact, the degrees of freedom are computed as follows, assuming that the population variances are equal.

Hence, it is found that the critical value for this two-tailed test is "t_c=3.25," for "\alpha=0.01" and "df=9."

The rejection region for this two-tailed test is "R=\{t:|t|>3.25\}."

Since it is assumed that the population variances are equal, the t-statistic is computed as follows:


"t={\bar{X_1}-\bar{X_2}\over \sqrt{\dfrac{(n_1-1)s_1^2+(n_2-1)s_2^2}{n_1+n_2-2}(\dfrac{1}{n_1}+\dfrac{1}{n_2})}}="

"={7.52-7.49\over \sqrt{\dfrac{(6-1)(0.024)^2+(5-1)(0.032)^2}{6+5-2}(\dfrac{1}{6}+\dfrac{1}{5})}}="

"=1.78"

Since it is observed that "|t|=1.78<3.35=t_c," it is then concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that the population mean "\mu_1" is different than  "\mu_2," at the 0.05 significance level.

Using the P-value approach: The p-value is "p=0.1089," and since "p=0.1089>0.01," it is concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that the population mean "\mu_1" is different than "\mu_2," at the 0.05 significance level.